# Sufficient and Necessary Barrier-like Conditions for Safety and Reach-avoid Verification of Stochastic Discrete-time Systems

### v2

@@ III Safety Verification
Since $v(\bm{x})$ satisfies (4), we have
$\begin{split}v(\bm{x})&\geq 1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x})+1_{\mathcal{X}}(\bm{x})\mathbb{E}_{\bm{\theta}}[v(\bm{f}(\bm{x},\bm{\theta}))]\\ &=1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\mathbb{E}_{\pi}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\\ &\geq 1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\mathbb{E}_{\pi}[1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(1))+1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(1))\mathbb{E}_{\bm{\theta}}[v(\bm{f}(\bm{\phi}_{\pi}^{\bm{x}}(1),\bm{\theta}))]]\\ &=\mathbb{E}_{\pi}[1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(1))]+1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\mathbb{E}_{\pi}[1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(1))v(\bm{\phi}_{\pi}^{\bm{x}}(2))]\\ &\geq\cdots\\ &\geq\mathbb{E}_{\pi}[1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+\sum_{i\in\mathbb{N}_{\geq 1}}\prod_{j=0}^{i-1}1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(i))]\\ &\geq V(\bm{x})\end{split}$
for $\bm{x}\in\mathbb{R}^{n}$. Therefore, according to Lemma 1, $\mathbb{P}_{\pi}(S_{\bm{x}_{0}})=1-V(\bm{x}_{0})\geq\epsilon_{1}$ holds.
@@
The resulting condition also serves as both a necessary and sufficient criterion for certifying lower bounds of safety probabilities. It is the one (9) in Proposition 3 in [29], which was derived using an auxiliary switched system and Ville’s Inequality [23]. In [29], only the sufficiency of the condition for safety verification was demonstrated. In addition, this condition serves as a typical example of the condition (3) with $\alpha=1$ and $\beta=0$ in Theorem 1 of [26], which investigates finite-time safety verification. It is important to note that while Proposition 2 in [19] also establishes a sufficient barrier-like condition for certifying upper bounds on the safety probability of avoiding unsafe sets when $\tilde{\alpha}=1$ and $\tilde{\beta}=0$, the safety probability pertains to a stopped process that stops evolving upon exiting the set $\mathcal{X}$. For interested readers, please refer to Proposition 2 in [19]. However, as discussed in Section I and in Remark 2, which is introduced later, the safety probability should be interpreted as the reach-avoid probability defined in Definition 4. $\blacksquare$
@@ IV Reach-avoid Verification
This section presents necessary and sufficient barrier-like conditions for the reach-avoid verification in Definition 4. Two cases are discussed in this section. The first case assumes that the system (1) will either leave the safe set $\mathcal{X}$ or enter the target set $\mathcal{X}_{r}$ in finite time almost surely. The second case considers the assumption that the specified lower bound $\epsilon_{2}$ is strictly less than the exact reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$, i.e., $\epsilon_{2}<\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$. These two cases are detailed in Subsection IV-A and IV-B, respectively.
@@ IV-A Reach-avoid Verification I
The subsection will formulate a necessary and sufficient barrier-like condition for the reach-avoid verification with the assumption that the system (1) will either leave the safe set $\mathcal{X}$ or enter the target set $\mathcal{X}_{r}$ in finite time almost surely. Like the one in Section III, this condition is also constructed via relaxing a Bellman equation. The Bellman equation is derived from a value function.
Let’s start with the value function $V(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$, which characterizes the exact reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}})$ for $\bm{x}\in\mathbb{R}^{n}$. We define the value function as follows:
@@
The value function $V(\bm{x})$ in (6) is equal to the reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}})$, i.e.,
$V(\bm{x})=\mathbb{P}_{\pi}(RA_{\bm{x}})$
for $\bm{x}\in\mathbb{R}^{n}$.
@@
Clearly, $\mathbb{E}_{\pi}[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))]=1_{\mathcal{X}_{r}}(\bm{x})$. In addition, since $\mathbb{E}_{\pi}[\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))]=\mathbb{P}_{\pi}(\wedge_{j=1}^{i-1}[\bm{\phi}_{\pi}^{\bm{x}}(j)\in\mathcal{X}\setminus\mathcal{X}_{r}]\wedge[\bm{\phi}_{\pi}^{\bm{x}}(i)\in\mathcal{X}_{r}])$ is the probability that the system (1) starting from $\bm{x}$ will enter the set $\mathcal{X}_{r}$ at time $t=i$ while staying within $\mathcal{X}\setminus\mathcal{X}_{r}$ before $i$. Thus, $\mathbb{P}_{\pi}(RA_{\bm{x}})=1_{\mathcal{X}_{r}}(\bm{x})+\sum_{i\in\mathbb{N}_{\geq 1}}\mathbb{P}_{\pi}(\wedge_{j=1}^{i-1}[\bm{\phi}_{\pi}^{\bm{x}}(j)\in\mathcal{X}\setminus\mathcal{X}_{r}]\wedge[\bm{\phi}_{\pi}^{\bm{x}}(i)\in\mathcal{X}_{r}])=V(\bm{x})$. ∎
We next will show that the value function (6) can be reduced to a solution to a Bellman equation via the dynamic programming principle. A controlled version of the Bellman equation can be found in [22].
@@
$\begin{split}V(\bm{x})=&1_{\mathcal{X}_{r}}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\mathbb{E}_{\pi}[1_{\mathcal{X}_{r}}(\bm{y})+\sum_{i\in\mathbb{N}_{\geq 1}}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{y}}(j))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{y}}(i))]\\ =&1_{\mathcal{X}_{r}}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\mathbb{E}_{\bm{\theta}}[1_{\mathcal{X}_{r}}(\bm{y})+\mathbb{E}_{\pi}[\sum_{i\in\mathbb{N}_{\geq 1}}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{y}}(j))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{y}}(i))]]\\ =&1_{\mathcal{X}_{r}}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\mathbb{E}_{\bm{\theta}}[V(\bm{f}(\bm{x},\bm{\theta}))]\\ \end{split}$
@@
Like the condition (4) in Theorem 1, we can also construct a necessary and sufficient condition for the safety verification scenario in [15], which is certifying upper bounds of the probability that the system eventually enters unsafe sets from an initial state while adhering to state-constrained sets, via relaxing the Bellman equation (7). It is shown in Proposition 3. The corresponding proof is shown in Appendix. In this proposition, $\mathcal{X}_{r}$ is a set of unsafe states and $\mathcal{X}$ is a state-constrained set. This condition is also a typical instance of condition (9) with $\alpha=1$ and $\beta=0$ in Theorem 3 in [26], which provides upper bounds of the reach-avoid probability in the finite-time reach-avoid verification.
@@
Let $\epsilon^{\prime}_{1}\in[0,1]$. There exists a function $v(\bm{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$ satisfying the following barrier-like condition:
@@
In addition, as discussed in Remark 1, a sufficient barrier-like condition is formulated in Proposition 2 with parameters $\tilde{\alpha}=1$ and $\tilde{\beta}=0$ in [19]. This condition can also be used for certifying upper bounds for the probability $\mathbb{P}_{\pi}(S^{\prime}_{\bm{x}_{0}})$. The primary distinction between this condition and the one presented in (8) is that the barrier function $B(\bm{x})$ in Proposition 2 in [19] does not require the condition $B(\bm{x})\geq 0$ for $\bm{x}\in\mathbb{R}^{n}\setminus\mathcal{X}$. $\blacksquare$
@@
When the system (1), starting from any state $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$, will leave the set $\mathcal{X}\setminus\mathcal{X}_{r}$ in finite time almost surely, i.e.,
$\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}_{\pi}^{\bm{x}}(k)\in\mathcal{X}\setminus\mathcal{X}_{r})=0,\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}.$
There are systems satisfying Assumption 1. For instance, this is the case when the safe set $\mathcal{X}$ is bounded and the system is subject to additive white Gaussian noise.
@@
Since $v(\bm{x})$ satisfies the Bellman equation (7), we have
$\begin{split}v(\bm{x})&=1_{\mathcal{X}_{r}}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\mathbb{E}_{\bm{\theta}}[v(\bm{f}(\bm{x},\bm{\theta}))]\\ &=1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\mathbb{E}_{\pi}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]\\ &=1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\mathbb{E}_{\pi}[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(1))+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(1))\mathbb{E}_{\bm{\theta}}[v(\bm{f}(\bm{\phi}_{\pi}^{\bm{x}}(1),\bm{\theta}))]]\\ &=\mathbb{E}_{\pi}[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(1))]+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))\mathbb{E}_{\pi}[1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(1))v(\bm{\phi}_{\pi}^{\bm{x}}(2))]\\ &=\cdots\\ &=\mathbb{E}_{\pi}[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))+\sum_{i\in\mathbb{N}_{\geq 1}}\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))]+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{i\rightarrow\infty}h_{i}(\bm{x})\\ &=V(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{i\rightarrow\infty}h_{i}(\bm{x})\end{split}$ (9)
for $\bm{x}\in\mathbb{R}^{n}$, where

### v3

@@ III Safety Verification
We first prove via induction that for all $k\in\mathbb{N}$,
$\begin{split}\zeta_{k}(\bm{x}):&=\mathbb{E}_{\pi}\left[\sum_{i=0}^{k}\prod_{j=0}^{i-1}1_{\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(j))\cdot 1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(i))\right]+\mathbb{E}_{\pi}\left[\prod_{j=0}^{k}1_{\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(j))\cdot v(\bm{\phi}^{\bm{x}}_{\pi}(k+1))\right]\leq v(\bm{x}).\end{split}$
Base Case ($k=0$):
$\begin{split}\zeta_{0}(\bm{x})&=\mathbb{E}_{\pi}\left[1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(0))\right]+\mathbb{E}_{\pi}\left[1_{\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(0))v(\bm{\phi}^{\bm{x}}_{\pi}(1))\right]\\ &=1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x})+1_{\mathcal{X}}(\bm{x})\mathbb{E}_{\bm{\theta}}[v(\bm{f}(\bm{x},\bm{\theta}))]\leq v(\bm{x}),\end{split}$
where the first equality follows from the convention that the empty product equals 1, and the inequality follows from condition (4).
Inductive Step: Assume $v(\bm{x})\geq\zeta_{k}(\bm{x})$ for some $k\geq 0$. Then:
$\zeta_{k+1}(\bm{x})$ $=\zeta_{k}(\bm{x})-\mathbb{E}_{\pi}\Bigg[\prod_{j=0}^{k}1_{\mathcal{X}}\big(\bm{\phi}^{\bm{x}}_{\pi}(j)\big)\,v\big(\bm{\phi}^{\bm{x}}_{\pi}(k+1)\big)\Bigg]$
$+\mathbb{E}_{\pi}\Bigg[\prod_{j=0}^{k}1_{\mathcal{X}}\big(\bm{\phi}^{\bm{x}}_{\pi}(j)\big)\Bigg(1_{\mathbb{R}^{n}\setminus\mathcal{X}}\big(\bm{\phi}^{\bm{x}}_{\pi}(k+1)\big)+1_{\mathcal{X}}\big(\bm{\phi}^{\bm{x}}_{\pi}(k+1)\big)\,\mathbb{E}_{\bm{\theta}}\bigg[v\big(\bm{\phi}^{\bm{x}}_{\pi}(k+2)\big)\bigg]\Bigg)\Bigg].$
Using condition (4) at state $\bm{\phi}^{\bm{x}}_{\pi}(k+1)$:
$\begin{split}v(\bm{\phi}^{\bm{x}}_{\pi}(k+1))\geq&1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(k+1))+1_{\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(k+1))\mathbb{E}_{\bm{\theta}}[v(\bm{\phi}^{\bm{x}}_{\pi}(k+2))],\end{split}$
we have $\zeta_{k+1}(\bm{x})\leq\zeta_{k}(\bm{x})\leq v(\bm{x})$.
By induction, $v(\bm{x})\geq\zeta_{k}(\bm{x})$ for all $k\in\mathbb{N}$. Since $\zeta_{k}(\bm{x})\geq 0$ for all $k\in\mathbb{N}$, $\lim_{k\to\infty}\zeta_{k}(\bm{x})$ exists. Taking $k\to\infty$, we have
$\begin{split}\lim_{k\to\infty}\zeta_{k}(\bm{x})&=\mathbb{E}_{\pi}\left[\sum_{i=0}^{\infty}\prod_{j=0}^{i-1}1_{\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(j))\cdot 1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(i))\right]+\lim_{k\to\infty}\mathbb{E}_{\pi}\left[\prod_{j=0}^{k}1_{\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(j))v(\bm{\phi}^{\bm{x}}_{\pi}(k+1))\right]\\ &\geq\mathbb{E}_{\pi}\left[\sum_{i=0}^{\infty}\prod_{j=0}^{i-1}1_{\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(j))\cdot 1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}^{\bm{x}}_{\pi}(i))\right]\\ &=V(\bm{x}).\end{split}$
Thus, $v(\bm{x})\geq V(\bm{x})$.
@@
The resulting condition also serves as both a necessary and sufficient criterion for certifying lower bounds of safety probabilities. It is the one (9) in Proposition 3 in [yu2023safe], which was derived using an auxiliary switched system and Ville’s Inequality [ville1939etude]. In [yu2023safe], only the sufficiency of the condition for safety verification was demonstrated. In addition, this condition serves as a typical example of the condition (3) with $\alpha=1$ and $\beta=0$ in Theorem 1 of [xue2024finite], which investigates finite-time safety verification. It is important to note that while Proposition 2 in [santoyo2021barrier] also establishes a sufficient barrier-like condition for certifying upper bounds on the safety probability of avoiding unsafe sets when $\tilde{\alpha}=1$ and $\tilde{\beta}=0$, the safety probability pertains to a stopped process that stops evolving upon exiting the set $\mathcal{X}$. For interested readers, please refer to Proposition 2 in [santoyo2021barrier]. However, as discussed in Section I and in Remark 2, which is introduced later, the safety probability should be interpreted as the reach-avoid probability defined in Definition 4. $\blacksquare$
@@ IV Reach-avoid Verification
This section presents necessary and sufficient barrier-like conditions for the reach-avoid verification in Definition 4. Two cases are discussed in this section. The first case assumes that, for every state in $\mathcal{X}\setminus\mathcal{X}_{r}$, the system (1) will either leave the safe set $\mathcal{X}$ or enter the target set $\mathcal{X}_{r}$ in finite time almost surely. The second case considers the assumption that the specified lower bound $\epsilon_{2}$ is strictly less than the exact reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$, i.e., $\epsilon_{2}<\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$. These two cases are detailed in Subsection IV-A and IV-B, respectively.
@@ IV-A Reach-avoid Verification I
This subsection formulates a necessary and sufficient barrier-like condition for reach-avoid verification, under the assumption that, for every state in $\mathcal{X}\setminus\mathcal{X}_{r}$, the system (1) will almost surely either enter the target set $\mathcal{X}_{r}$ or exit the safe set $\mathcal{X}$ in finite time. Similar to the one in Section III, this condition is also constructed by relaxing a Bellman equation. The Bellman equation is derived from a value function.
We begin by introducing the value function $V(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$, which characterizes the exact reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}})$ for $\bm{x}\in\mathbb{R}^{n}$. We define the value function as follows:
@@
The value function $V(\bm{x})$ in (6) is equal to the reach-avoid probability $\mathbb{P}_{\pi}(RA_{\bm{x}})$, i.e., $V(\bm{x})=\mathbb{P}_{\pi}(RA_{\bm{x}})$ for $\bm{x}\in\mathbb{R}^{n}$.
@@
By definition, $\mathbb{E}_{\pi}[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))]=1_{\mathcal{X}_{r}}(\bm{x})$. In addition, since $\mathbb{E}_{\pi}[\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))]=\mathbb{P}_{\pi}(\wedge_{j=0}^{i-1}[\bm{\phi}_{\pi}^{\bm{x}}(j)\in\mathcal{X}\setminus\mathcal{X}_{r}]\wedge[\bm{\phi}_{\pi}^{\bm{x}}(i)\in\mathcal{X}_{r}])$ is the probability that the system (1) starting from $\bm{x}$ will enter the set $\mathcal{X}_{r}$ at time $t=i$ while staying within $\mathcal{X}\setminus\mathcal{X}_{r}$ before $i$, where $i\geq 1$. Thus, we have
$\begin{split}&\mathbb{E}_{\pi}[1_{\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(0))]+\sum_{i\in\mathbb{N}_{\geq 1}}\mathbb{E}_{\pi}[\prod_{j=0}^{i-1}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathbb{R}^{n}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(i))]\\ &=\mathbb{P}_{\pi}(\bm{\phi}_{\pi}^{\bm{x}}(0)\in\mathcal{X}_{r})+\sum_{i\in\mathbb{N}_{\geq 1}}\mathbb{P}_{\pi}(\wedge_{j=0}^{i-1}[\bm{\phi}_{\pi}^{\bm{x}}(j)\in\mathcal{X}\setminus\mathcal{X}_{r}]\wedge[\bm{\phi}_{\pi}^{\bm{x}}(i)\in\mathcal{X}_{r}])\\ &=\mathbb{P}_{\pi}(RA_{\bm{x}}).\end{split}$
Consequently, $\mathbb{P}_{\pi}(RA_{\bm{x}})=V(\bm{x}).$
We next will show that the value function (6) can be reduced to a solution to a Bellman equation via the dynamic programming principle. A controlled version of the Bellman equation can be found in [summers2010verification].
@@
[…]
@@
Similar to the condition (4) in Theorem 1, we can also construct a necessary and sufficient condition for the safety verification scenario in [prajna2007framework], which is certifying upper bounds of the probability that the system eventually enters unsafe sets from an initial state while adhering to state-constrained sets, by relaxing the Bellman equation (7). It is shown in Proposition 3. The proof is shown in Appendix. In this proposition, $\mathcal{X}_{r}$ is a set of unsafe states and $\mathcal{X}$ is a state-constrained set. This condition is also a typical instance of condition (9) with $\alpha=1$ and $\beta=0$ in Theorem 3 in [xue2024finite], which provides upper bounds of the reach-avoid probability in the finite-time reach-avoid verification.
@@
Let $\epsilon^{\prime}_{1}\in[0,1]$. There exists a function $v(\bm{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$ satisfying the barrier-like condition:
@@
In addition, as discussed in Remark 1, a sufficient barrier-like condition is formulated in Proposition 2 with parameters $\tilde{\alpha}=1$ and $\tilde{\beta}=0$ in [santoyo2021barrier]. This condition can also be used for certifying upper bounds for the probability $\mathbb{P}_{\pi}(S^{\prime}_{\bm{x}_{0}})$. The primary distinction between this condition and the one presented in (8) is that the barrier function $B(\bm{x})$ in Proposition 2 in [santoyo2021barrier] does not require the condition $B(\bm{x})\geq 0$ for $\bm{x}\in\mathbb{R}^{n}\setminus\mathcal{X}$. $\blacksquare$
@@
For every initial state $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$, the system (1) exits the set $\mathcal{X}\setminus\mathcal{X}_{r}$ in finite time almost surely; that is, $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}_{\pi}^{\bm{x}}(k)\in\mathcal{X}\setminus\mathcal{X}_{r})=0,\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$.
There are systems satisfying Assumption 1. For instance, consider the stochastic system $\bm{x}(k+1)=\bm{g}(\bm{x}(k))+\bm{\theta}(k)$, where $\bm{\theta}(k)$ is a i.i.d. Gaussian disturbance. Let the initial state $\bm{x}(0)$ lie within a bounded set $\mathcal{X}\setminus\mathcal{X}_{r}$. Since the additive noise has unbounded support, there is a non-zero probability that the trajectory will eventually exit any bounded set. In fact, with probability one, the trajectory will leave $\mathcal{X}\setminus\mathcal{X}_{r}$ in finite time. Therefore, the condition $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}_{\pi}^{\bm{x}}(k)\in\mathcal{X}\setminus\mathcal{X}_{r})=0,\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$ holds, satisfying Assumption 1.
@@
We first show that $v(\bm{x})=V(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{k\to\infty}h_{k}(\bm{x})$ for $\bm{x}\in\mathbb{R}^{n}$, where
$h_{k}(\bm{x}):=\mathbb{E}_{\pi}\left[\prod_{j=1}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))\cdot v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\right]$
for $k\in\mathbb{N}$. We note that when $k=0$, the product is taken over an empty index set, and by convention, the empty product equals 1. Thus, $h_{0}(\bm{x})=\mathbb{E}_{\pi}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]=\mathbb{E}_{\bm{\theta}}[v(\bm{f}(\bm{x},\bm{\theta}))]$.
For this sake, we prove by induction that for all $k\in\mathbb{N}$,